Solve the following system of equations.
x - 2y = 5
2x - 4y = 10
Thank you!

Answer by stanbon(75887)

Solve the following system of equations.
x - 2y = 5
2x - 4y = 10
-----
Multiply thru the 1st equation by 2 to get:
2x - 4y = 10
2x - 4y = 10
-----
The system is dependent (one equation is a multiple of the other).
The solution is all ordered pairs of the form (x, (x-5)/2)
-----------------
Cheers,
Stan H.
=====================